One-half of Heather's age two years from now plus one-third of her age three years ago is twenty years. How old is she now?


1
Expert's answer
2021-11-05T18:19:30-0400

"\frac {x+2}2+\frac{x-3}3=20,"

"3x+6+2x-6=120,"

x=24 (years).